OpenStudy (pucusana):

Use the linear equation 3(x+1)=3x+3 to answer parts (a)–(d). Does x=5 satisfy the equation above? Explain. Is x=-8 a solution of the equation above? Explain. Is x=1/2 a solution of the equation above? Explain. What interesting fact about the equation 3(x+1)=3x+3 is illuminated by the answers to parts (a), (b), and (c)? Why do you think this is true?

OpenStudy (pucusana):

anyone?

OpenStudy (pucusana):

help me?

OpenStudy (pucusana):

please?

OpenStudy (anonymous):

a) Yes. 3(5+1)=3(6)=18 ; 3(5)+3=15+3=18 b) Yes. 3(-8+1)=3(-7)=-21 ; 3(-8)+3=-24+3=-21 c) Yes. 3(1/2+1)=3(3/2)=9/2 ; 3(1/2)+3=(3/2)+(6/2)=9/2 d) I'm not entirely sure what they mean by interesting fact, that's pretty vague. Maybe that if you distribute the 3(x+1) you get 3x+3.